# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ 3. The topological invariant
If two time-reversal symmetric projection-valued maps $P_{0},P_{1}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ are unitarily equivalent or homotopic, then $\delta(P_{0})=\delta(P_{1})$.
@@
Suppose that there is a unitary-valued map $U:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ such that $P_{1}(k)=U(k)P_{0}(k)U(k)^{-1}$ with TU(k)=U(k)TTU(k)=U(-k)T. Then we can use Theorem 4.2 to decompose $P_{0}=P_{0}^{-}(k)\oplus P_{0}^{+}(k)$ with $P_{0}^{\pm}:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ and $TP_{0}^{+}(k)=P_{0}^{-}(-k)T$. Clearly
$P_{1}(k)=U(k)P_{0}^{-}(k)U(k)^{-1}\oplus U(k)P_{0}^{+}(k)U(k)^{-1}$
is a decomposition of $P_{1}(k)$ such that $TP_{1}^{+}(k)=P_{1}^{-}(-k)T$. Since the Chern number is stable under unitary equivalences (Proposition 2.5), we have that $\operatorname{Ch}(P_{0}^{-})=\operatorname{Ch}(P_{1}^{-})$. So, it follows that:
@@
Instead, suppose that there is an homotopy $P_{t}(k)$ with $TP_{t}(k)=P_{t}(-k)T$ for $t\in[0,1]$ that connects $P_{0}$ with $P_{1}$. then we can divide $[0,1]$ into a partition $t_{j},t_{j+1}]_{j\in\{0,\cdots,J-1\}}$ such that $t_{0}=0,t_{J}=1$ and such that $\|P_{t}(k)-P_{t_{j}}(k)\|<1$ for all $t\in[t_{j},t_{j+1}],k\in\mathbb{T}^{2}$. Then we can use the Kato-Nagy formula 7 to define the unitary-valued maps
$U_{j}(k)=P_{t_{j+1}}(k)P_{t_{j}}(k)+(\mathds{1}-P_{t_{j+1}}(k))(\mathds{1}-P_{t_{j}}(k))\left[\mathds{1}-(P_{t_{j}}(k)-P_{t_{j+1}}(k))^{2}\right]^{-1/2}$
for all $j\in\{0,\cdots,J-1\}$. It is easy to see that $U_{j}(k)$ is a symmetric unitary equivalence between $P_{t_{j}}(k)$ and $P_{t_{j+1}}(k)$ for all $j\in\{0,\cdots,J-1\}$, so, using the previous point, we see that $\delta$ is preserved at each step and therefore $\delta(P_{0})=\delta(P_{1})$. ∎
4. Splitting of time-reversal symmetric projections
In this section, we want to describe how it is possible to express the invariant $\delta$ as the parity of a Chern number. To do so, we need to use a specific formulation of the Chern number and we also need to give an algorithm that allows us to decompose $P$ as a pair $P=P^{+}+P^{-}$.
Remark 4.1.
As stated in [panati2007triviality, monaco2015symmetry], if $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ is a time-reversal symmetric projection-valued map, then $\operatorname{Ch}(P)=0$.
Indeed, we can use Proposition 2.3 to define a unitary-valued map $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ with $P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}$ and $TU(t,k_{2})=U(-t,-k_{2})T$. Then if $\{v_{j}\}_{j\in\{1,\cdots,n\}}$ is a quaternionic orthonormal basis of $\operatorname{Ran}(P(0,0))$ and $v_{j+n}=Tv_{j}$ for $j\in\{1,\cdots,n\}$ as in Lemma A.1, we will have the unitary matrix $\alpha\in U(2n)$ with:
$\begin{split}[\alpha]_{a,b}(k_{2})&=\left\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\right\rangle=\left\langle U(-\pi,k_{2})v_{a},U(\pi,k_{2})v_{b}\right\rangle=\\ &=\left\langle TU(\pi,k_{2})v_{b},TU(-\pi,k_{2})v_{a}\right\rangle=\left\langle U(-\pi,-k_{2})v_{b},U(\pi,-k_{2})v_{a}\right\rangle=\\ &=[\alpha]_{b,a}(-k_{2}).\end{split}$
This means that $\det(\alpha(k_{2}))=\det(\alpha(-k_{2}))$, but a map with this property has a winding number equal to zero due to Lemma A.4.
@@ 4. Splitting of time-reversal symmetric projections
Given a time-reversal symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ and an integer $h\in\mathbb{Z}$, if $e^{i\pi h}=\delta(P)$, it is always possible to decompose $P$ as $P(k)=P^{+}(k)\oplus P^{-}(k)$ with $\operatorname{Ch}(P^{-})=h$. Vice versa, if $P$ can be decomposed as $P(k)=P^{+}(k)\oplus P^{-}(k)$, then $e^{i\pi\operatorname{Ch}(P^{-})}=\delta(P)$.
Proof.
We want to start using the same notation used in Definition 3.1, which means that $P(t,k_{2})$ is spanned by $\{u_{j}(t,k_{2})\}_{j\in\{1,\cdots,2n\}}$ with constraints:
$Tu_{j}(t,k_{2})=\sum_{a=1}^{2n}[-J]_{j,a}u_{a}(-t,-k_{2}),\quad u_{j}(-\pi,k_{2})=\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u_{a}(\pi,k_{2})$
where $\alpha$ is a periodic unitary-valued map such that $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$ and $J=\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}$.
Equivalent conditions.  Anticipating some results from the next section, we notice that the existence of a decomposition is tantamount to the existence of a symmetric frame with appropriate pseudo-periodic conditions. In fact, given a decomposition $P(k)=P^{+}(k)\oplus P^{-}(k)$, we can apply Theorem 5.1 to $P(k_{1},k_{2})$ to construct a frame $\{v_{j}(k_{1},k_{2})\}_{j\in\{1,\cdots,n\}}$ of $P^{-}(k_{1},k_{2})$ with complete periodicity except for the first vector, that instead satisfies
$v_{1}(\pi,k_{2})=e^{i\operatorname{Ch}(P^{-})k_{2}}v_{1}(-\pi,k_{2})=e^{ihk_{2}}v_{1}(-\pi,k_{2}).$
Then
$\{v_{n+j}(t,k_{2})=-Tv_{j}(-t,-k_{2})\}_{j\in\{1,\cdots,n\}}$
is a frame of $P^{+}(t,k_{2})=T^{-1}P^{-}(-t,-k_{2})T$ with complete periodicity except for the first vector that satisfies
$v_{n+1}(\pi,k_{2})=e^{-i\operatorname{Ch}(P^{-})k_{2}}v_{n+1}(-\pi,k_{2})=e^{-ihk_{2}}v_{n+1}(-\pi,k_{2})\,.$
The union of the two collections of vectors will constitute a symmetric frame for $P$ and this also implies that $\operatorname{Ch}(P^{+})=-\operatorname{Ch}(P^{-})$. Vice versa, if we have a frame with these properties, we can take $P^{-}(k)$ as the span of the first $n$ vectors and $P^{+}(k)$ as the span of the remaining $n$ vectors.
At the same time, the existence of a frame of this type is tantamount to the existence of a unitary-valued map $\beta:[-\pi,\pi]\times\mathbb{T}^{1}\to U(2n)$ such that $v_{j}(t,k_{2})=\sum_{a=1}^{2n}[\beta(t,k_{2})]_{j,a}u_{a}(t,k_{2})$. The frame is symmetric if and only if $J\beta(t,k_{2})=\overline{\beta(-t,-k_{2})}J$ because:
Instead, the frame satisfies the pseudoperiodic condition if and only if
$\beta(\pi,k_{2})=\Lambda(k_{2})\beta(-\pi,k_{2})\alpha(k_{2})$
where $\Lambda(k)$ is a diagonal matrix with
$\Lambda(k)=\operatorname{diag}(e^{ihk_{2}},\underbrace{1,\cdots,1}_{n-1},e^{-ihk_{2}},\underbrace{1,\cdots,1}_{n-1}).$
This is true because:
$\begin{split}v_{j}(\pi,k_{2})&=\sum_{a=1}^{2n}[\Lambda(k_{2})]_{j,a}v_{a}(-\pi,k_{2})\Leftrightarrow\\ \sum_{a=1}^{2n}[\beta(\pi,k_{2})]_{j,a}u_{a}(\pi,k_{2})&=\sum_{a=1}^{2n}[\Lambda(k_{2})\beta(-\pi,k_{2})]_{j,a}u_{a}(-\pi,k_{2})\Leftrightarrow\\ \sum_{a=1}^{2n}[\beta(\pi,k_{2})]_{j,a}u_{a}(\pi,k_{2})&=\sum_{a=1}^{2n}[\Lambda(k_{2})\beta(-\pi,k_{2})\alpha(k_{2})]_{j,a}u_{a}(\pi,k_{2})\end{split}$
However, the symmetry condition forces a constraint on $\beta$: in fact if $J\beta(0,k_{2})=\overline{\beta(0,-k_{2})}J$, then $\beta(0,0),\beta(0,\pi)\in\mathrm{Sp}(2n)$. This implies that $\det(\beta(0,0))=\det(\beta(0,\pi))=1$. Moreover, $\det(\beta(0,k_{2}))=\overline{\det(\beta(0,-k_{2})}$, so we can use the Lemma A.3 to conclude that the winding number in the second direction $\det(\beta(t,\cdot))]$ must be even.
Decomposition when $\delta(P)=1$. Now we want to investigate under which assumption it is possible to find such a $\beta$ when $h$ is even. First of all, we can look at the points $(\pi,0),(\pi,\pi)$ in the torus. There it must be true that:
$\beta(\pi,0)=\Lambda(0)\beta(-\pi,0)\alpha(0),\quad\beta(\pi,\pi)=\Lambda(\pi)\beta(-\pi,\pi)\alpha(\pi).$
However, if $h\in 2\mathbb{Z}$, then $\Lambda(0)=\Lambda(\pi)=\mathbb{I}_{2n}$. If we use the periodicity in $k_{2}$ and the symmetry constraint, we get that $\beta$ must satisfy
$\beta(\pi,k_{2})=\Lambda(k_{2})J\overline{\beta(\pi,-k_{2})}J^{-1}\alpha(k_{2})$
and in particular
$J\beta(\pi,0)^{t}J^{-1}\beta(\pi,0)=\alpha(0),\quad J\beta(\pi,\pi)^{t}J^{-1}\beta(\pi,\pi)=\alpha(\pi).$
This means that, since $\det(\Lambda(k_{2}))\equiv 1$, it is true that
$\begin{cases}\det(\beta(\pi,k_{2}))=\overline{\det(\beta(\pi,-k_{2}))}\det(\alpha(k_{2}))\\ \det(\beta(\pi,0))^{2}=\det(\alpha(0))\\ \det(\beta(\pi,\pi))^{2}=\det(\alpha(\pi))\end{cases}$
Now we can choose a function $\lambda:[0,\pi]\to\mathbb{R}$ such that $\det(\beta(\pi,k_{2}))=e^{i\lambda(k_{2})}$, as well as a function $\mu:[-\pi,\pi]\to\mathbb{R}$ such that $\det(\alpha(k_{2}))=e^{i\mu(k_{2})}$ and with $\mu(0)=2\lambda(0)$. This means that:
$\det(\beta(\pi,k_{2}))=\begin{cases}e^{i\lambda(k_{2})}&\mbox{for }k_{2}\in[0,\pi]\\ e^{-i\lambda(-k_{2})+i\mu(-k_{2})}&\mbox{for }k_{2}\in[-\pi,0]\end{cases}$
So we can write $\det(\beta(\pi,k_{2}))=e^{i\phi(k_{2})}$ where $\phi:[-\pi,\pi]\to\mathbb{R}$ is a continuous but not periodic function such that:
$\phi(k_{2})=\begin{cases}\lambda(k_{2})&\mbox{for }k_{2}\in[0,\pi]\\ \lambda(-k_{2})+\mu(-k_{2})&\mbox{for }k_{2}\in[-\pi,0]\end{cases}$
This implies that
$\det(\beta(\pi,\cdot))]=\frac{\lambda(\pi)-[-\lambda(\pi)+\mu(\pi)]}{2\pi}=\frac{2\lambda(\pi)-\mu(\pi)}{2\pi}$
So, if we interpret $\beta(\pi,0)$ as $\gamma_{0}$ and $\beta(\pi,\pi)$ as $\gamma_{\pi}$, then we can say that $\lambda(0)=\lambda_{0},\lambda(\pi)=\lambda_{\pi}$ and obtain that
$e^{i\pi[\det(\pi,k_{2})]}=e^{i\pi\frac{2\lambda(\pi)-\mu(\pi)}{2\pi}}=\delta(P)$
This means that if $\delta(P)$ is odd, it is impossible to find $\beta$ because $\delta$ does not depend on the choices of $\beta(\pi,0)$ and $\beta(\pi,\pi)$ (Proposition 3.2) and the symmetry constraint imposed that $\det(\beta(t,\cdot))]$ must be even for every $t$. Instead if $\delta(P)=2h$ is even we can use the Sylvester theorem to find two self-adjoint operators $L,R$ with $JLJ^{-1}=L^{t},JRJ^{-1}=R^{t}$ and such that $\alpha(0)=e^{iL},\alpha(\pi)=e^{iR}$, and then define:
$\beta(\pi,k_{2})=\begin{cases}\exp\left[i\frac{\pi-k_{2}}{2\pi}L+i\frac{k_{2}}{2\pi}R\right]&\mbox{for }k_{2}\in[0,\pi]\\ \Lambda(k_{2})J\overline{\beta(\pi,-k_{2})}J^{-1}\alpha(k_{2})&\mbox{for }k_{2}\in[-\pi,0]\end{cases}$
This is continuous and periodic in $k_{2}$ because when $k_{2}=0$ we get:
$e^{iL/2}=\Lambda(0)Je^{-i\overline{L}/2}J^{-1}e^{iL}=e^{-iJL^{t}J^{-1}/2}e^{iL}=e^{iL/2}$
and similar computations hold for $k_{2}=\pi$. The previous argument states that $\det(\beta(\pi,k_{2}))]$ is even equal to $2r$. Then we can also define
$\beta(0,k_{2})=\operatorname{diag}(e^{irk_{2}},\underbrace{1,\cdots,1}_{n-1},e^{irk_{2}},\underbrace{1,\cdots,1}_{n-1})$
and it is trivial to check that $\beta(0,k_{2})$ satisfies the symmetry constraint and has $\det(\beta(0,k_{2}))]=2r=[\det(\beta(\pi,k_{2}))]$. Finally we can use Theorem A.7 to find a homotopy $\beta_{t}(k_{2})$ in $U(2n)$ with $\beta_{0}(k_{2})=\beta(0,k_{2}),\beta_{1}(k_{2})=\beta(\pi,k_{2})$ and then define:
$\beta(t,k_{2})=\begin{cases}\beta_{t/\pi}(k_{2})&\mbox{for }t\in[0,\pi]\\ J^{-1}\overline{\beta_{-t/\pi}(-k_{2})}J&\mbox{for }t\in[-\pi,0]\end{cases}$
This satisfies all the desired conditions, so if $\delta(P)$ is even, we can build a decomposition with $\operatorname{Ch}(P^{-})=h$ for any even number $h$.
Decomposition when $\delta(P)=-1$.  Now we want to investigate under which assumption it is possible to find a decomposition with $\operatorname{Ch}(P^{-})$ odd. It is immediate to check that the conditions are identical to the even case with a small but crucial difference. If $\operatorname{Ch}(P^{-})$ is even, it holds that:
$\Lambda(0)=\mathbb{I}_{2n},\quad\Lambda(\pi)=\begin{pmatrix}-1&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&-1&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}$
Once again, if we use the periodicity in $k_{2}$ and the symmetry constraint, we get that $\beta$ must satisfy
$\beta(\pi,k_{2})=\Lambda(k_{2})J\overline{\beta(\pi,-k_{2})}J^{-1}\alpha(k_{2})$
and in particular
$J\beta(\pi,0)^{t}J^{-1}\beta(\pi,0)=\alpha(0),\quad J\beta(\pi,\pi)^{t}J^{-1}\Lambda(\pi)\beta(\pi,\pi)=\alpha(\pi).$
With some computation, it is easy to check that
$\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}\begin{pmatrix}-1&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&-1&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}$=
$\begin{pmatrix}i&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&i&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}^{t}\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}\begin{pmatrix}i&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&i&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}$
meaning that if
$Y=\operatorname{diag}(i,\underbrace{1,\cdots,1}_{n-1},i,\underbrace{1,\cdots,1}_{n-1})$
we have that $\det(Y)=-1$ and the last condition can be rewritten as:
$J(Y\beta(\pi,\pi))^{t}J^{-1}(Y\beta(\pi,\pi))=\alpha(\pi).$
We can replicate the previous step of the proof to show that it is possible to define $\beta$ if and only if $\frac{2\lambda(\pi)-\mu(\pi)}{2\pi}\in 2\mathbb{Z}.$ In this case we can interpret $\beta(\pi,0)$ as $\gamma_{0}$ and $Y\beta(\pi,\pi)$ as $\gamma_{\pi}$. Then we can impose $\lambda_{0}=-i\ln(\det(\pi,0))=\lambda(0)$, so that $\mu$ does not need to be changed, and $\lambda_{\pi}=-i\ln(\det(Y\beta(\pi,\pi)))=\lambda(\pi)+\pi$ because
$\det(Y\beta(\pi,\pi))=-\det(\beta(\pi,\pi))=e^{i(\lambda(\pi)+\pi)}.$
We conclude that, in this case, we have that:
$\exp\left(i\frac{2\lambda(\pi)-\mu(\pi)}{2}\right)=\exp\left(i\frac{2\lambda_{\pi}-\mu(\pi)-\pi}{2}\right)=-\delta(P)$
Since we need an even winding number of $\det(\beta(\pi,k_{2}))]$, if $\delta(P)=-1$ it is impossible to construct a frame with $h$ odd, instead if $\delta(P)=1$ we can proceed exactly as before and obtain the thesis. Namely, it is possible to decompose $P=P^{+}\oplus P^{-}$ with $\operatorname{Ch}(P^{-})$ odd if and only if $\delta(P)$ is odd. ∎

### v2

@@ 5.2. Addivity and stability of the $\mathbb{Z}_{2}$ invariant
If two time-reversal symmetric projection-valued maps $P_{0},P_{1}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ are Murray–von Neumann equivalent, unitarily equivalent or homotopic (in the sense of definition 1.4), then $\delta(P_{0})=\delta(P_{1})$.
@@
Suppose that there is a Murray–von Neumann equivalence $V$ or a unitary equivalence $U$. Through Theorem 5.1 we can obtain symmetric splitting $P_{0}=P_{0}^{-}(k)+P_{0}^{+}(k)$ with $P_{0}^{\pm}:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ and $TP_{0}^{+}(k)=P_{0}^{-}(-k)T$. Clearly
$P_{1}(k)=V(k)P_{0}^{-}(k)V(k)^{*}+V(k)P_{0}^{+}(k)V(k)^{*}$
or
$P_{1}(k)=U(k)P_{0}^{-}(k)U(k)^{-1}+U(k)P_{0}^{+}(k)U(k)^{-1}$
are valid splittings of $P_{1}(k)$ in the sense of Definition 1.7. Since the Chern number is stable under Murray–von Neumann and unitary equivalences (Proposition 3.3), we have that $\operatorname{Ch}(P_{0}^{-})=\operatorname{Ch}(P_{1}^{-})$. So, it follows that:
@@
Instead, suppose that there is an homotopy $P_{t}(k)$ with $TP_{t}(k)=P_{t}(-k)T$ for $t\in[0,1]$ that connects $P_{0}$ with $P_{1}$. Then we can divide $[0,1]$ into a partition $t_{j},t_{j+1}]$, $j\in\{0,\cdots,J-1\}$}, such that $t_{0}=0,t_{J}=1$ and such that $\|P_{t}(k)-P_{t_{j}}(k)\|<1$ for all $t\in[t_{j},t_{j+1}]$, $k\in\mathbb{T}^{2}$. Then we can use the Kato-Nagy formula (7) to define the unitary-valued maps $U_{j}(k):=U_{P_{t_{j}}(k)\to P_{t_{j+1}}(k)}$ for all $j\in\{0,\cdots,J-1\}$. It is easy to see that $U_{j}(k)$ is a symmetric unitary equivalence between $P_{t_{j}}(k)$ and $P_{t_{j+1}}(k)$ for all $j\in\{0,\cdots,J-1\}$, so, using the previous point, we see that $\delta$ is preserved at each step and therefore $\delta(P_{0})=\delta(P_{1})$. ∎
Lemma 5.3.
Consider two time-reversal symmetric projection-valued maps $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$, $Q:\mathbb{T}^{2}\to\operatorname{Proj}_{2m}(\mathcal{H})$ such that $P(k)Q(k)\equiv 0$: then $\delta(P)\delta(Q)=\delta(P+Q)$.
Proof.
We can use Theorem 5.1 to find two symmetric splittings $P(k)=P^{+}(k)+P^{-}(k)$, $Q(k)=Q^{+}(k)+Q^{-}(k)$. Then it is obvious that
$P(k)+Q(k)=[P^{-}(k)+Q^{-}(k)]+[P^{+}(k)+Q^{+}(k)]$
defines a symmetric splitting of $P+Q$. From Lemma 3.4 we have $\operatorname{Ch}(P^{-}+Q^{-})=\operatorname{Ch}(P^{-})+\operatorname{Ch}(Q^{-})$, so we can conclude using 5.1 that
$\delta(P+Q)=e^{i\pi\operatorname{Ch}(P^{-}+Q^{-})}=e^{i\pi\operatorname{Ch}(P^{-})}\cdot e^{i\pi\operatorname{Ch}(Q^{-})}=\delta(P)\cdot\delta(Q).\qed$
5.3. From pseudo-periodic frames to periodic Parseval frames
Theorem 5.1 yields the following conclusion: every continuous time-reversal symmetric projection-valued map $P$ splits as a sum of two orthogonal continuous time-reversal symmetric maps
$P(k)=P_{2}(k)+P_{2(n-1)}(k)\,,\quad P_{2}(k)\,P_{2(n-1)}(k)\equiv 0\,$,
where $P_{2}(k)$ is the orthogonal projection on the $2$-dimensional subspace spanned by the Kramer pair of vectors in a pseudo-periodic symmetric frame which are not periodic, and $P_{2(n-1)}(k)=P(k)-P_{2}(k)$ is the projection onto the (2n2)(2n-2)-dimensional subspace spanned by the periodic vectors in the frame. Moreover, by the same result, we have that $\delta(P)=\delta(P_{2})\in\mathbb{Z}_{2}$, and that $\delta(P_{2(n-1)})=0\in\mathbb{Z}_{2}$ (as $P_{2(n-1)}$ admits a continuous, periodic and symmetric frame).
Consider now the projection-valued map
$Q(k):=P_{2}(k)\oplus P_{2}(k)\quad\text{on }\mathcal{H}\oplus\mathcal{H}\,.$
It is clear that $Q(k)$ is a rank-4 projection-valued map which is continuous, periodic, and time-reversal symmetric with respect to the time-reversal symmetry operator $T\oplus T$ on $\mathcal{H}\oplus\mathcal{H}$. Moreover, by additivity of the $\mathbb{Z}_{2}$ invariant (Lemma 5.3), we have
$\delta(Q)=\delta(P_{2})\,\delta(P_{2})=1\bmod 2$
and therefore, by Theorem 5.1, it admits a periodic symmetric frame
$\{u_{1}(k),u_{2}(k),u_{3}(k)=Tu_{1}(-k),u_{4}(k)=Tu_{2}(-k)\}.$
Call $w_{j}(k):=\Pi(u_{j}(k))$, $1\leq j\leq 4$, where $\Pi\colon\mathcal{H}\oplus\mathcal{H}\to\mathcal{H}$ is the projection on the first summand in the direct sum. While not orthonormal, the vectors $\{w_{1}(k),\ldots,w_{4}(k)\}$ are still continuous, periodic, time-reversal symmetric, and span $\operatorname{Ran}P_{2}(k)$; moreover, the Parseval condition holds, namely
$\psi=\sum_{j=1}^{4}\left\langle w_{j}(k),\psi\right\rangle\,w_{j}(k)\quad\text{for all }\psi\in\operatorname{Ran}P_{2}(k)\,.$
Following the lines of [cornean2019parseval] (compare also [auckley2018parseval]) we have thus proved that
@@ 5.3. From pseudo-periodic frames to periodic Parseval frames
Any continuous, periodic and time-reversal symmetric rank-$2n$ projection-valued map can be spanned by a Parseval frame consisting of $2n+2$ continuous, periodic and symmetric vectors, namely of $n+1$ continuous and periodic Kramers pairs.
5.4. Emerging spin–type simmetry
Let $P$ be a time–reversal symmetric projection–valued map, and consider a continuous unitary–valued map $S:\mathbb{T}^{2}\to U(\mathcal{H})$ that squares to the identity, commutes with $P$, and satisfies the following compatibility relation with the time-reversal symmetry $T$:
$TS(k)=-S(-k)T,\qquad\forall k\in\mathbb{T}^{2}.$
We will refer to such a map as a spin-type symmetry of $P$. Given a spin-type symmetry $S$, the Spin-Chern number of $P$ is defined as the integer-valued quantity
$\mathrm{Ch}_{S}(P):=\frac{\operatorname{Ch}\!\left(P\frac{S+\mathds{1}}{2}\right)-\operatorname{Ch}\!\left(P\frac{S-\mathds{1}}{2}\right)}{2}.$
Indeed, since $S^{2}=\mathds{1}$, the operators $(S\pm\mathds{1})/2$ define continuous projection-valued maps. Moreover, since $[S,P]=0$, the operators $P^{\pm}:=P(S\pm\mathds{1})/2$ are themselves projection-valued, and hence their Chern numbers $\operatorname{Ch}(P^{\pm})$ are well defined. Since $(S(k)\pm\mathds{1})/2$ are the spectral projections onto the eigenspaces of $S(k)$, the decomposition $P=P^{+}+P^{-}$ is orthogonal. By Lemma 3.4, one then has
$0=\operatorname{Ch}(P)=\operatorname{Ch}(P^{+})+\operatorname{Ch}(P^{-}),$
which implies that $\operatorname{Ch}(P^{+})$ and $\operatorname{Ch}(P^{-})$ coincide modulo $2$, so that their half-difference $\operatorname{Ch}_{S}(P)$ is an integer. Furthermore, the same relation yields
(11) $\operatorname{Ch}_{S}(P)=\operatorname{Ch}(P^{\pm})=\operatorname{Ch}\bigg(P\frac{S\pm\mathds{1}}{2}\bigg)\mod 2.$
Note that we have not assumed that $S$ is independent of $k$, although this is typically the case in physical models where the Spin-Chern number is defined. In fact, the operator $S$ usually corresponds to a conserved spin component and acts as an internal symmetry of the system, that is, it acts only on the second factor of the tensor product $L^{2}(X)\otimes\mathbb{C}^{n}$ on which the Hamiltonian is defined. When such an operator is decomposed through the Bloch–Floquet–Zak transform, its fiber representation is always independent of the quasi-momentum variable $k$, since the transform acts only on the first factor of the tensor product.
Clearly, the independence of $S$ from $k$ is not required in order to define the associated Spin-Chern number. In what follows, we show that for any time-reversal symmetric projection-valued map $P$ it is always possible to construct a spin-type symmetry $S$, and that the $\mathbb{Z}_{2}$ invariant of $P$ can be computed through the associated Spin-Chern number as
$\delta(P)=(-1)^{\operatorname{Ch}_{S}(P)}.$
In general, such an operator $S$ will be $k$-dependent and not unique; however, the Spin-Chern number of $P$ depends on the choice of $S$ only modulo $2$.
It is worth noting that the idea of computing $\mathbb{Z}_{2}$-valued invariants via additional symmetries has appeared previously in the literature. In [kellendonk2019cycliccohomologygradedcralgebras], a highly general framework based on Van Daele $K$-theory and cyclic cohomology is developed, where torsion-valued pairings give rise to $\mathbb{Z}_{2}$ topological invariants. In Section 7.3 of that work, this general theory is applied to tight-binding, possibly non-periodic systems subject to fermionic time-reversal symmetry. It is shown that any such system admits, up to stabilization, a representative of its $K$-theory class carrying an additional spin-type symmetry, and that the resulting torsion-valued pairing, which expresses the expected $\mathbb{Z}_{2}$ invariant, reduces to its Spin-Chern number associated with the additional symmetry. The content of this section may therefore be viewed as a direct, non-$K$-theoretic analogue of these results in the periodic, not necessarily tight-binding, setting. In particular, we emphasize that the construction of the spin-type symmetry presented below requires no stabilization of the projection-valued map $P$.
Before proving the main theorem, we introduce the following technical lemma, which essentially allows one to construct symmetric, continuous, and periodic frames for infinite-dimensional time-reversal symmetric projection-valued maps.